OpenStudy (anonymous):

Suppose that the celsius temperature at the point (x,y) in the xy-plane is T(x,y)=xsin(2y) and that distance in the xy-plane is measured in meters. A particle is moving clockwise around the circle of radius 1 m centered at the origin at a constant rate of 2m/sec. A) how fast is the temp experienced by the particle changing in degress cel per meter at the point: P(1/2, sqrt(3)/2) * I don't need the answer just how to find the unit vector, which I know is

OpenStudy (anonymous):

Are you studying parametric equations?

OpenStudy (anonymous):

multivar calc

OpenStudy (anonymous):

so, parametric eq are fair game

OpenStudy (anonymous):

not sure what unit vector you're talking about, but it's pretty quick: express the circle in terms of x(t) and y(t). plug those into T(x,y) take the derivative with respect to time plug in the time that corresponds to (1/2, sqrt(3)/2)

OpenStudy (anonymous):



OpenStudy (anonymous):

I guess maybe you're talking about the direction of velocity...?

OpenStudy (anonymous):

I'm trying to solve it using gradients, I apologize, this is probably very obvious, but my ability to visualize these types of problems is very limited

OpenStudy (anonymous):

I have a solutions manual, and everything makes sense except for the very first step in which it says simply: "The unit tangent vector @ (1/2, Sqrt(3)/2) in the direction of motion is: u= (sqrt(3)/2i-1/2j)" I just can't see how they calculated that.

OpenStudy (anonymous):

if you're asking about the direction of velocity at the point (1/2, sqrt(3)/2) then plug those coord.s into the equations of motion I gave: 1/2 = cos(2*t) sqrt(3) /2 = sin(-2*t) find t find the time derivative of x(t), y(t) (-2sin(2*t) , -2cos(2*t) ) plug in the t you found to find the velocity vector at that point :)

OpenStudy (anonymous):

that's how they calculated it.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

sure, don't forget that the vector you get will need to be divided by magnitude to get a unit vector... (the vector you get is just <Vx, Vy> )

OpenStudy (anonymous):

Got it, thanks again.